# On Marginal Stability in Low Temperature Spherical Spin Glasses

### v2

@@ 1.5 Technical Preliminaries
The next proposition provides smoothness estimates for $H_{N}$, ensuring for example that the radial derivative has typical order $O(1)$, while the spherical gradient has typical norm $O(\sqrt{N})$. The operator norm of a tensor $\boldsymbol{A}\in({\mathbb{R}}^{N})^{\otimes k}$ is
${\|\boldsymbol{A}\|}_{\rm op}=\max_{\|{\boldsymbol{\sigma}}^{1}\|_{2},\ldots,\|{\boldsymbol{\sigma}}^{k}\|_{2}\leq 1}|\langle\boldsymbol{A},{\boldsymbol{\sigma}}^{1}\otimes\cdots\otimes{\boldsymbol{\sigma}}^{k}\rangle|\,.$
We denote by ${\mathscr{H}}_{N}\simeq\bigoplus_{p\geq 1}{\mathbb{R}}^{N^{p}}$ the set of mixed $p$-spin Hamiltonians on ${\mathcal{S}}_{N}$, which for fixed $\xi$ can be identified with the coefficients ${\boldsymbol{G}}^{(p)}$.
@@
[HS24b, Proposition 1.1] shows that (1.17) and (1.18) hold with probability $1-e^{-cN}$. The same holds for (1.19) by the Borell-TIS inequality. ∎
@@
$|H_{N}({\boldsymbol{\sigma}})/N-{\mathrm{GS}}_{N}(\xi)|$ $\leq\varepsilon,$ (1.20)
@@
The first bound (1.20) follows by (1.19). This implies (1.21) and (1.22) because if one of them did not hold, then for $H_{N}\in K_{N}$, Taylor expanding $H_{N}$ near ${\boldsymbol{\sigma}}$ would show (1.20) is false for some $\varepsilon^{\prime}$. ∎
The next standard lemma shows that $\partial_{{\mathrm{rad}}}H_{N}({\boldsymbol{\sigma}})$ determines the bulk spectral edge of $\nabla_{\mathrm{sp}}^{2}H_{N}({\boldsymbol{\sigma}})$ uniformly over ${\boldsymbol{\sigma}}\in{\mathcal{S}}_{N}$. In particular, (1.9) and (1.8) are equivalent (the case $N(1-\eta)\leq j\leq N-K$ follows by negating $H_{N}$, which preserves its law). We note that in Section 3, it is important to take $j$ constant in Lemma 1.11.
Lemma 1.11 ([Sub21, Lemma 3]).
@@
$\big{|}{\boldsymbol{\lambda}}_{j}\big{(}\nabla_{\mathrm{sp}}^{2}H_{N}({\boldsymbol{\sigma}})\big{)}-2\xi^{\prime\prime}(1)-\partial_{{\mathrm{rad}}}H_{N}({\boldsymbol{\sigma}})\big{|}\leq\varepsilon$
@@
Assume $H_{N}\in K_{N}$. Then $A_{\beta^{-1/2}}$ (recall (1.7)) satisfies $\mu_{\beta}(A_{\beta^{-1/2}})\geq 1-e^{-cN}$ for $\beta\geq\beta_{*}(\xi)$ large.
@@ 1.6 Characterization of the Minimizer in the Parisi Formula
Since the Parisi functional ${\mathcal{Q}}$ is strictly convex, the unique minimizer $(\zeta,L)$ is characterized by first-order stationarity conditions. These will be useful later, and we review them now. Given $(\zeta,L)\in{\mathcal{K}}$, define
@@
The characterization below is primarily from [CS17, Theorem 2]. The fact that $T$ consists of finitely many intervals is proved in [JT18, Corollary 1.6] at finite temperature, and the proof is essentially the same. One combines [JT17, Theorem 1.13] and the observation that by analyticity, $\big{(}\frac{1}{\sqrt{\xi^{\prime\prime}}}\big{)}^{\prime\prime}$ changes sign finitely many times on $[0,1]$.
@@
There exists a unique $(L,\zeta)$ attaining the infimum (1.4), which is characterized by the following. Here ${\rm supp}(\zeta)\subseteq[0,1)$ denotes the set of points of increase of $\zeta$, i.e. the support of $\nu$.
$G(1)=0;\quad\quad\min_{q\in[0,1]}g(q)=0;\quad\quad T\subseteq{\rm supp}(\zeta).$
@@
If $1$ is not an isolated point in $T$, then $\xi$ exhibits full RSB endpoint behavior.
@@
By the last assertion of Proposition 1.13, $T$ must contain a non-trivial interval $(1-\varepsilon(\xi),1)$. Then $g^{\prime\prime}(1)=0$ by definition of $G$, which implies (1.10). ∎
@@ 2.1 Radial Derivative at Approximate Ground States
We fix $\xi$ and $\eta<0.01$ with $\xi(1+\eta)<\infty$, and let $\xi_{t}(q)=\xi(t^{2}q)$ for $t\in(1-\eta,1+\eta)$. To explain this definition, note that by (1.1) or (1.2), the function on ${\mathcal{S}}_{N}$ defined by ${\boldsymbol{\sigma}}\to H_{N}(t{\boldsymbol{\sigma}})$ is precisely a spherical spin glass with mixture $\xi_{t}$.
@@
be the corresponding minimizers in the zero-temperature Parisi formula.
@@
There exists $C=C(\xi)>0$ such that for all $t\in(1-\eta,1+\eta)$,
@@
We treat $C=C(\xi)$ as a constant that may vary from line to line. In the topologically trivial case that $\zeta_{t}$ is identically 00, we easily find $L_{t}=\sqrt{\xi_{t}^{\prime}(1)}$ and so $\hat{\zeta}_{t}(1)=1/\sqrt{\xi_{t}^{\prime}(1)}$, so the conclusion is obvious. We assume this is not the case below. Noting that $\hat{\zeta}_{t}(\cdot)$ is a decreasing function of $t$, it follows that $\hat{\zeta}_{t}(1/2)\leq C$ is bounded independently of $t$ since
$2{\mathrm{GS}}(\xi_{t})$ $\geq 2{\mathcal{Q}}(\zeta_{t},L_{t};\xi_{t})\geq\int_{0}^{1}\xi_{t}^{\prime\prime}(q)\hat{\zeta}(q){\rm d}q\geq\int_{0}^{1/2}\xi_{t}^{\prime\prime}(q)\hat{\zeta}(q){\rm d}q$
$\geq\frac{\hat{\zeta}_{t}(1/2)}{2}\int_{0}^{1/2}\xi_{t}^{\prime\prime}(q)\leavevmode\nobreak\ {\rm d}q\geq\frac{\hat{\zeta}_{t}(1/2)}{2}\big{(}\xi^{\prime}(1/2)-\xi^{\prime}(0)\big{)}\geq\Omega(\hat{\zeta}_{t}(1/2)).$
Therefore $\hat{\zeta}_{t}(q)\geq L_{t}-C$ for all $q\in[1/4,1/2]$. Since $\xi_{t}^{\prime\prime}(1/4)\geq 1/C$ for all $t\in[1-\eta,1+\eta]$, we find that $L_{t}\leq C$ is also bounded independently of $t$.
We now turn to the second estimate. Let $A=\xi_{1+\eta}^{\prime\prime}(1)^{-1/2}\leq\xi_{t}^{\prime\prime}(1)^{-1/2}$, and suppose without loss of generality that $\hat{\zeta}_{t}(1)<A$. Let $q_{*}\in[0,1)$ satisfy $\hat{\zeta}_{t}(q_{*})=A$, and consider
$\zeta(q)=\begin{cases}\zeta(q),\quad q\leq q_{*};\\ \zeta(q_{*}),\quad q\geq q_{*}.\end{cases}$
Then by definition $\hat{\zeta}(q)\in[\hat{\zeta}_{t}(q),A]$ for all $q\in[0,1]$, so if $\zeta\neq\zeta_{t}$, we easily find that ${\mathcal{Q}}(\zeta,L_{t})<{\mathcal{Q}}(\zeta_{t},L_{t})$ which is a contradiction. (By similar reasoning, if $\hat{\zeta}(0)<A$, then increasing $L_{t}$ would decrease ${\mathcal{Q}}$, so $q_{*}$ actually exists.) Therefore $\zeta_{t}(q)=z$ must be constant on $q\in[q_{*},1)$.
@@
Next, let $q^{*}\leq q_{*}$ be the largest point in the support of $\zeta_{t}$, with $q^{*}=0$ if $\zeta_{t}$ is identically 00. Recalling the notation from (1.23), we have $G(1)=G(q^{*})$ which means
@@
$\xi_{t}^{\prime}(1)-\xi_{t}^{\prime}(q^{*})\geq\frac{1}{2z\hat{\zeta}_{t}(1)}\implies\hat{\zeta}_{t}(1)\geq\frac{1}{(1-q^{*})zC^{\prime}(\xi)}.$
However by our assumptions and the first part above, $(1-q^{*})z=\hat{\zeta}_{t}(1)-\hat{\zeta}_{t}(q^{*})\leq L_{t}\leq C$. Thus $\hat{\zeta}_{t}(1)$ is bounded away from 00 depending only on $\xi$ in all cases. ∎
@@
$\frac{{\rm d}}{{\rm d}s}{\mathrm{GS}}(\xi_{s})=\frac{{\rm d}}{{\rm d}t}{\mathcal{Q}}(\zeta_{s},L_{s};\xi_{t})\big{|}_{t=s}=\xi_{s}^{\prime}(0)L_{s}+\int_{0}^{1}\big{(}2\xi_{s}^{\prime\prime}(q)+q\xi_{s}^{\prime\prime\prime}(q)\big{)}\hat{\zeta}_{s}(q)\leavevmode\nobreak\ {\rm d}q.$ (2.2)
@@
Lemma 2.1 shows ${\mathrm{GS}}(\xi_{s})$ is defined as an infimum over the compact set ${\mathcal{K}}_{C}\subseteq{\mathcal{K}}$; Proposition 1.13 ensures the minimizer is unique. This easily implies $s\mapsto(\zeta_{s},L_{t})$ is continuous. The first equality now follows from the envelope theorem [MS02, Theorem 2 and Corollary 4], assuming said derivative of ${\mathcal{Q}}$ exists. Indeed dominated convergence shows it is given by the explicit formula above. (As $C$ is fixed, there are no analytic issues for $q\approx 1$.) ∎

### v3

@@ 1.5 Technical Preliminaries
The next proposition provides smoothness estimates for $H_{N}$, ensuring for example that the radial derivative has typical order $O(1)$, while the spherical gradient has typical norm $O(\sqrt{N})$. The operator norm of a tensor $\boldsymbol{A}\in({\mathbb{R}}^{N})^{\otimes p}$ is
${\|\boldsymbol{A}\|}_{\rm op}=\max_{\|{\boldsymbol{\sigma}}^{1}\|_{2},\ldots,\|{\boldsymbol{\sigma}}^{p}\|_{2}\leq 1}|\langle\boldsymbol{A},{\boldsymbol{\sigma}}^{1}\otimes\cdots\otimes{\boldsymbol{\sigma}}^{p}\rangle|\,.$
We denote by ${\mathscr{H}}_{N}\simeq\bigoplus_{p\geq 1}{\mathbb{R}}^{N^{p}}$ the set of mixed $p$-spin Hamiltonians on ${\mathcal{S}}_{N}$, which for fixed $\xi$ can be identified with the coefficient tensors ${\boldsymbol{G}}^{(p)}\in({\mathbb{R}}^{N})^{\otimes p}$
@@
[HS25b, Proposition 2.3] shows that (1.18) and (1.19) hold with probability $1-e^{-cN}$. The same holds for (1.20) by the Borell-TIS inequality. ∎
The following description of the near-global maxima is elementary. We explain it in detail, in part as a warm-up for the proof of Theorem 1.3 in Subsections 2.2 and 2.3, which ends with a similar spherical Taylor expansion.
@@
$|H_{N}({\boldsymbol{\sigma}})/N-{\mathrm{GS}}(\xi)|$ $\leq\varepsilon,$ (1.21)
@@
The first bound (1.21) follows by (1.20). To deduce (1.22) and (1.23), we will show that if one of them does not hold, then for $H_{N}\in K_{N}$, then Taylor expanding $H_{N}$ near ${\boldsymbol{\sigma}}$ would falsify (1.20). In both cases, it will be convenient to use the following spherical Taylor expansion, which is easily verified using Proposition 1.9 and justifies the definition of the Riemannian Hessian $\nabla_{\mathrm{sp}}^{2}$. For ${\boldsymbol{v}}\in{\mathcal{S}}_{N}\cap{\boldsymbol{\sigma}}^{\perp}$, we have $\cos(\varepsilon){\boldsymbol{\sigma}}+\sin(\varepsilon){\boldsymbol{v}}\in{\mathcal{S}}_{N}$ and:
$H_{N}(\cos(\varepsilon){\boldsymbol{\sigma}}+\sin(\varepsilon){\boldsymbol{v}})$ $=H_{N}({\boldsymbol{\sigma}})+\varepsilon\langle\nabla_{\mathrm{sp}}H_{N}({\boldsymbol{\sigma}}),{\boldsymbol{v}}\rangle+\frac{\varepsilon^{2}}{2}\langle\nabla_{\mathrm{sp}}^{2}H_{N}({\boldsymbol{\sigma}}),{\boldsymbol{v}}^{\otimes 2}\rangle+O(\varepsilon^{3}N)$ (1.24)
$=H_{N}({\boldsymbol{\sigma}})+\varepsilon\langle\nabla_{\mathrm{sp}}H_{N}({\boldsymbol{\sigma}}),{\boldsymbol{v}}\rangle+O(\varepsilon^{2}N).$
Here the hidden implicit constants are uniform in ${\boldsymbol{\sigma}}$ and ${\boldsymbol{v}}$ and $\varepsilon$, for all $H_{N}\in K_{N}$.
For (1.22), let us assume $H_{N}\in K_{N}$ and (without loss of generality) that $\varepsilon$ is small depending on $\xi$. Suppose that $\delta<\varepsilon^{3}/4$ is such that (1.20) holds with constant $\varepsilon^{3}/4$ (noting that $\varepsilon>0$ was arbitrary in the original statement and hence can be adjusted separately in the three estimates to be proved). We will show that (1.22) then follows with constant $\varepsilon$. Indeed if it does not, we let ${\boldsymbol{v}}=\sqrt{N}\nabla_{\mathrm{sp}}H_{N}({\boldsymbol{\sigma}})/\|\nabla_{\mathrm{sp}}H_{N}({\boldsymbol{\sigma}})\|$ and $\varepsilon^{\prime}=\varepsilon\|\nabla_{\mathrm{sp}}H_{N}({\boldsymbol{\sigma}})\|/\sqrt{N}$. Then
$H_{N}(\cos(\varepsilon^{\prime}){\boldsymbol{\sigma}}+\sin(\varepsilon^{\prime}){\boldsymbol{v}})$ $=H_{N}({\boldsymbol{\sigma}})+\varepsilon\|\nabla_{\mathrm{sp}}H_{N}({\boldsymbol{\sigma}})\|^{2}+O(\varepsilon^{\prime 2}N)$
$\geq H_{N}({\boldsymbol{\sigma}})+\frac{\varepsilon\|\nabla_{\mathrm{sp}}H_{N}({\boldsymbol{\sigma}})\|^{2}}{2}$
$\geq H_{N}({\boldsymbol{\sigma}})+\frac{\varepsilon^{3}N}{2}$
$>N\Big{(}{\mathrm{GS}}(\xi)+\frac{\varepsilon^{3}}{4}\Big{)}.$
Here the second inequality follows from the assumption that $\|\nabla_{\mathrm{sp}}H_{N}({\boldsymbol{\sigma}})\|\geq\varepsilon\sqrt{N}$ and we used ${\boldsymbol{\sigma}}\in A_{\delta}$ for $\delta<\varepsilon^{3}/4$ to obtain the last inequality. Comparing the initial and final expressions above contradicts our assumption that (1.20) holds with constant $\varepsilon^{3}/4$.
For the last assertion, we again suppose $H_{N}\in K_{N}$ and $\varepsilon$ is small depending on $\xi$. Here we assume (1.23) does not hold with constant $\varepsilon$ but that both (1.22) and (1.20) hold with constant $\varepsilon^{5}/8$. Then let ${\boldsymbol{w}}/\sqrt{N}$ be the top unit eigenvector of $\nabla_{\mathrm{sp}}^{2}H_{N}({\boldsymbol{\sigma}})$, viewed as a quadratic form on ${\boldsymbol{\sigma}}^{\perp}$. With $\varepsilon^{\prime\prime}=\varepsilon{\boldsymbol{\lambda}}_{1}$ for ${\boldsymbol{\lambda}}_{1}={\boldsymbol{\lambda}}_{1}(\nabla_{\mathrm{sp}}^{2}H_{N}({\boldsymbol{\sigma}}))>\varepsilon$ we find
$H_{N}(\cos(\varepsilon^{\prime\prime}){\boldsymbol{\sigma}}+\sin(\varepsilon^{\prime\prime}){\boldsymbol{w}})$ $=H_{N}({\boldsymbol{\sigma}})+O(\varepsilon^{\prime\prime}\|\nabla_{\mathrm{sp}}H_{N}({\boldsymbol{\sigma}})\|\sqrt{N})+\frac{(\varepsilon^{\prime\prime})^{2}{\boldsymbol{\lambda}}_{1}N}{2}+O((\varepsilon^{\prime\prime})^{3}N)$
$\geq H_{N}({\boldsymbol{\sigma}})+(\varepsilon^{\prime\prime})^{2}{\boldsymbol{\lambda}}_{1}N/4$
$\geq H_{N}({\boldsymbol{\sigma}})+\varepsilon^{5}N/4$
$>N\Big{(}{\mathrm{GS}}(\xi)+\frac{\varepsilon^{5}}{8}\Big{)}.$
Again we used ${\boldsymbol{\sigma}}\in A_{\delta}$ (now with $\delta<\varepsilon^{5}/8$) to obtain the last inequality, and the final inequality contradicts the assumption that (1.20) holds (with constant $\varepsilon^{5}/8$). ∎
The next standard lemma states that $\partial_{{\mathrm{rad}}}H_{N}({\boldsymbol{\sigma}})$ determines the bulk spectrum of $\nabla_{\mathrm{sp}}^{2}H_{N}({\boldsymbol{\sigma}})$ uniformly over ${\boldsymbol{\sigma}}\in{\mathcal{S}}_{N}$. In particular, (1.10) and (1.9) are equivalent (the case $N(1-\eta)\leq j\leq N-K$ follows by negating $H_{N}$, which preserves its law). We note that in Section 3, it is important to take $j$ constant in Lemma 1.11.
Lemma 1.11 ([Sub21, Lemma 1.3]).
@@
$\big{|}{\boldsymbol{\lambda}}_{j}\big{(}\nabla_{\mathrm{sp}}^{2}H_{N}({\boldsymbol{\sigma}})\big{)}-2\xi^{\prime\prime}(1)^{{1/2}}{+}\partial_{{\mathrm{rad}}}H_{N}({\boldsymbol{\sigma}})\big{|}\leq\varepsilon$
@@
Assume $H_{N}\in K_{N}$. Then for $\beta\geq\beta_{*}(\xi)$ large, $A_{\beta^{-1/2}}$ (recall (1.8)) satisfies $\mu_{\beta}(A_{\beta^{-1/2}})\geq 1-e^{-cN}$.
@@ 1.6 Characterization of the Minimizer in the Parisi Formula
Since the Parisi functional ${\mathcal{Q}}$ is strictly convex, the unique minimizer $(\zeta,L)$ is characterized by first-order stationarity conditions. These will be useful later, and we review them now. Given $(\zeta,L)\in{\mathcal{K}}$ (recall (1.4)), define
@@
The characterization below is primarily from [CS17, Theorem 2]. In the finite temperature case [JT18, Corollary 1.6] proves that $T$ consists of finitely many intervals, and the zero temperature case is identical. (One combines [JT17, Theorem 1.13] and the observation that by analyticity, $\big{(}\frac{1}{\sqrt{\xi^{\prime\prime}}}\big{)}^{\prime\prime}$ changes sign finitely many times on $[0,1]$.)
@@
The unique $(L,\zeta)$ attaining the infimum (1.5) is characterized as follows, where ${\rm supp}(\zeta)\subseteq[0,1)$ denotes the set of points of increase of $\zeta$ (i.e. the support of $\nu$ in (1.26)):
$G(1)=0;\quad\quad\min_{q\in[0,1]}g(q)=0;\quad\quad T{\supseteq}{\rm supp}(\zeta).$
@@
Suppose $1$ is not an isolated point in $T$. Then $\xi$ exhibits full RSB endpoint behavior. If $\xi$ is not a quadratic polynomial, then $\zeta$ is moreover strictly increasing in a neighborhood of $1$.
@@
Proposition 1.13 implies that $T$ contains a non-trivial interval $(a,1)$. On this interval, we must have $g^{\prime\prime}(q)=0$, which easily rearranges to
This yields the first assertion. For the latter, note that the right-hand side is an analytic function on an open complex neighborhood of $[a,1]$ and is not identically zero since $\xi$ (being generic) is not quadratic. Thus it has finitely many zeros. As $\zeta$ is increasing by definition, this implies $\zeta$ is strictly increasing on $(1-\varepsilon_{*}(\xi),1)$ as desired. ∎
Remark 1.15.
The converse of the first assertion in Corollary 1.14 need not hold. Indeed, suppose the non-trivial interval $[t,s]\subseteq(0,1)$ is a connected component of the set $T$ (see [AZ22b] for explicit constructions of such behavior). As in the proof of Corollary 1.14, one then finds $\hat{\zeta}(t)=\xi^{\prime}(t)^{-1/2}$. It is not hard to show that $\xi_{\sqrt{t}}(q)=\xi(tq)$ has full-RSB endpoint behavior at $1$, but that $1$ is an isolated point of $T(\xi_{\sqrt{t}})$. This is because the associated Parisi order parameter is a reparametrization of $\zeta$ (see e.g. [Sub24, Proposition 11] or [HS23a, Proposition 4.3]). We expect such counterexamples to essentially comprise the boundary of the full RSB phase.
Remark 1.16.
As previously mentioned, a large portion of Corollary 1.5 follows from the weaker assumption that strict full RSB endpoint behavior is absent (i.e. that $1$ is isolated in $T$), without requiring the approach in Theorem 1.3. We assume $(a,1)$ is disjoint from $T$, and again that $\xi$ is even. It follows from two-replica interpolation bounds at zero temperature (see e.g. [AC18b, Theorem 6 and 11], or [CGPR19, Theorem 3] as summarized in [Sel24, Lemma 6.4]) that for small $\delta$ depending on $\varepsilon$, with probability $1-e^{-cN}$ there does not exist any pair $({\boldsymbol{\sigma}},{\boldsymbol{\sigma}}^{\prime})\in A_{\delta}\times A_{\delta}$ such that $R({\boldsymbol{\sigma}},{\boldsymbol{\sigma}}^{\prime})\in(a+\varepsilon,1-\varepsilon)$. Taking $\varepsilon$ small depending on $a$, we may define a graph on vertex set $A_{\delta}$ such that ${\boldsymbol{\sigma}},{\boldsymbol{\sigma}}^{\prime}$ are adjacent if $R({\boldsymbol{\sigma}},{\boldsymbol{\sigma}}^{\prime})\geq 1-\varepsilon$. Then the property above implies all connected components of this graph have Euclidean diameter at most $2\sqrt{\varepsilon N}$, and are pairwise $10\sqrt{\varepsilon N}$-separated.
To our knowledge these components need not be connected with respect to the usual topology on ${\mathcal{S}}_{N}$, and indeed could have many critical points for $H_{N}$. Still, the existence of such a clustering suffices for our proofs of Corollaries 1.6 and 1.7 to go through unchanged. We thank the anonymous referee for helpful observations that led to this remark.
@@ 2.1 Radial Derivative at Approximate Ground States
We fix $\xi$ and $\eta<0.01$ with $\xi(1+{2\eta})<\infty$, and let $\xi_{t}(q)=\xi(t^{2}q)$ for $t\in(1-\eta,1+\eta)$. To explain this definition, note that by (1.2), the function on ${\mathcal{S}}_{N}$ defined by ${\boldsymbol{\sigma}}\to H_{N}(t{\boldsymbol{\sigma}})$ is precisely a spherical spin glass with mixture $\xi_{t}$.
@@
be the corresponding (unique) minimizers in the zero-temperature Parisi formula.
@@
There exists $C=C(\xi,\eta)>0$ such that for all $t\in(1-\eta,1+\eta)$,
@@
We treat $C=C(\xi,\eta)$ as a constant that may vary from line to line. In the “topologically trivial” case that $\zeta_{t}$ is identically 00, we find ${L_{t}=\hat{\zeta}_{t}(1)=1/\sqrt{\xi_{t}^{\prime}(1)}}$, so the conclusion is obvious. We assume this is not the case below. Noting that $\hat{\zeta}_{t}(\cdot)$ is a decreasing function for each $t$, it follows that $\hat{\zeta}_{t}(1/2)\leq C$ is bounded independently of $t$ since
${2\,{\mathrm{GS}}(\xi_{1+\eta})}\geq 2\,{\mathrm{GS}}(\xi_{t})$ $\geq 2{\mathcal{Q}}(\zeta_{t},L_{t};\xi_{t})$
$\geq\int_{0}^{1}\xi_{t}^{\prime\prime}(q){\hat{\zeta}_{t}(q)}{\rm d}q$
$\geq\int_{0}^{1/2}\xi_{t}^{\prime\prime}(q){\hat{\zeta}_{t}(q)}{\rm d}q$
$\geq\frac{\hat{\zeta}_{t}(1/2)}{2}\int_{0}^{1/2}\xi_{t}^{\prime\prime}(q)\leavevmode\nobreak\ {\rm d}q$
$\geq\frac{\hat{\zeta}_{t}(1/2)}{2}{\big{(}\xi_{t}^{\prime}(1/2)-\xi_{t}^{\prime}(0)\big{)}}$
$\geq\Omega(\hat{\zeta}_{t}(1/2)).$
(Since $\xi(1+2\eta)<\infty$, the coefficients of $\xi_{1+\eta}$ decay exponentially and so ${\mathrm{GS}}(\xi_{1+\eta})<\infty$. We also used the easy fact that $GS(\xi_{t})$ is increasing in $t$ in the first step.) Next we note that $\hat{\zeta}_{t}(\cdot)$ is concave and non-negative on $[0,1]$ since $\zeta_{t}(\cdot)$ is increasing. Therefore
$\hat{\zeta}_{t}(0)\leq 2\hat{\zeta}_{t}(1/2)-\hat{\zeta}_{t}(1)\leq 2\hat{\zeta}_{t}(1/2)\leq 2C$
is bounded depending only on $\xi$. Since $L_{t}=\hat{\zeta}_{t}(0)$ we conclude the first estimate.
We now turn to the second estimate. Let
$A=\frac{1}{2\sqrt{{\mathrm{GS}}(\xi_{1+\eta})^{2}+\xi_{1+\eta}^{\prime\prime}(1)}}\leq\frac{1}{2\sqrt{{\mathrm{GS}}(\xi_{t})^{2}+\xi_{t}^{\prime\prime}(1)}}\leq\frac{1}{2{\mathrm{GS}}(\xi_{t})}$
and assume that $\hat{\zeta}_{t}(1)<A$ (else we are already done). We first claim that $\hat{\zeta}_{t}(0)\geq A$. Indeed if not, then $\hat{\zeta}_{t}(q)<\frac{1}{2{\mathrm{GS}}(\xi_{t})}$ for all $q\in[0,1]$ and so
${\mathcal{Q}}(\zeta_{t},L_{t};\xi_{t})\geq\frac{1}{2}\int_{0}^{1}\frac{{\rm d}q}{\hat{\zeta}_{t}(q)}>{\mathrm{GS}}(\xi_{t})$
which is impossible. Given that $\hat{\zeta}_{t}(0)\geq A>\hat{\zeta}_{t}(1)$, it follows by continuity that there exists $q_{*}\in[0,1)$ such that $\hat{\zeta}_{t}(q_{*})=A$. We will show that $\zeta_{t}$ is constant on $[q_{*},1]$. Indeed consider the modified order parameter
${\zeta^{*}_{t}}(q)=\begin{cases}\zeta(q),\quad q\leq q_{*};\\ \zeta(q_{*}),\quad q\geq q_{*}.\end{cases}$
We aim to show that $\zeta^{*}_{t}(q)=\zeta_{t}(q)$ for all $q\in[0,1]$, via the inequality ${\mathcal{Q}}(\zeta_{t}^{*},L_{t};\xi_{t})\leq{\mathcal{Q}}(\zeta_{t},L_{t};\xi_{t})$. Setting
$\hat{\zeta}^{*}_{t}(q)\equiv L_{t}-\int_{0}^{q}\zeta^{*}_{t}(u){\rm d}u,$
by construction $\hat{\zeta}_{t}^{*}$ is decreasing with $\hat{\zeta}_{t}^{*}(q_{*})=A$, and so $\hat{\zeta}_{t}^{*}(q)\leq A$ for all $q\geq q_{*}$. Further $\zeta^{*}_{t}(q)\leq\zeta_{t}(q)$ for each $q\in[0,1]$, which implies by integration that $\hat{\zeta}^{*}_{t}(q)\geq\hat{\zeta}_{t}(q)$ for all $q$. Combining, we have for $q\geq q_{*}$ that
$\hat{\zeta}_{t}^{*}(q)\in[\hat{\zeta}_{t}(q),A]\subseteq[\hat{\zeta}_{t}(q),\xi_{t}^{\prime\prime}(q)^{-1/2}]$
which by monotonicity of $x\mapsto\xi_{t}^{\prime\prime}(q)x+\frac{1}{x}$ on $x\in(0,\xi_{t}^{\prime\prime}(q)^{-1/2})$ implies
$\xi_{t}^{\prime\prime}(q)\hat{\zeta}_{t}(q)+\frac{1}{\hat{\zeta}_{t}(q)}\geq\xi_{t}^{\prime\prime}(q)\hat{\zeta}^{*}_{t}(q)+\frac{1}{\hat{\zeta}^{*}_{t}(q)}.$ (2.2)
For $q\leq q_{*}$, we also have (2.2), in fact with equality. Inspecting the definition of ${\mathcal{Q}}$, it then follows that
${\mathcal{Q}}(\zeta_{t},L_{t};\xi_{t})\geq{\mathcal{Q}}(\zeta^{*}_{t},L_{t};\xi_{t}).$
By minimality of $(\zeta_{t},L_{t})$ we must have equality, which means equality holds in (2.2) for every $q$. Hence $\zeta_{t}=\zeta^{*}_{t}$ and so $\zeta_{t}(q)$ must be constant on $q\in[q_{*},1)$. Below we let $z=\zeta_{t}(q_{*})$ be this constant value.
@@
Next, let $q^{*}\in[0,q_{*}]$ be the largest point in the support of $\zeta_{t}$ (we already ruled out the case that $\zeta_{t}$ is identically zero). Recalling the notation from (1.25), we have $G(1)=G(q^{*})$ which means
@@
$\xi_{t}^{\prime}(1)-\xi_{t}^{\prime}(q^{*})\geq\frac{1}{2z\hat{\zeta}_{t}(1)}\implies\hat{\zeta}_{t}(1)\geq\frac{1}{(1-q^{*})zC^{\prime}(\xi)},$
where
$C^{\prime}(\xi)=2\xi_{1+\eta}^{\prime\prime}(1)\geq 2\xi_{t}^{\prime\prime}(1)\geq 2\cdot\frac{\xi_{t}^{\prime}(1)-\xi_{t}^{\prime}(q^{*})}{1-q_{*}}.$
However by our assumptions and the first part above,
[…]
Thus $\hat{\zeta}_{t}(1)$ is bounded away from 00 depending only on $\xi$ in all cases. ∎
@@
${\frac{1}{s}}\frac{{\rm d}}{{\rm d}s}{\mathrm{GS}}(\xi_{s})={\frac{1}{s}}\left(\frac{{\rm d}}{{\rm d}t}{\mathcal{Q}}(\zeta_{s},L_{s};\xi_{t})\big{|}_{t=s}\right)=\xi_{s}^{\prime}(0)L_{s}+\int_{0}^{1}\big{(}2\xi_{s}^{\prime\prime}(q)+q\xi_{s}^{\prime\prime\prime}(q)\big{)}\hat{\zeta}_{s}(q)\leavevmode\nobreak\ {\rm d}q.$ (2.3)
@@
Lemma 2.1 shows ${\mathrm{GS}}(\xi_{s})$ is defined as an infimum over the compact set ${\mathcal{K}}_{C}\subseteq{\mathcal{K}}$; Proposition 1.13 ensures the minimizer is unique. This easily implies $s\mapsto(\zeta_{s},{L_{s}})$ is continuous. Using the just-established compactness, the first equality now follows from the envelope theorem [MS02, Theorem 2 and Corollary 4], assuming said derivative of ${\mathcal{Q}}$ exists. Indeed dominated convergence and the chain rule yields the explicit formula above. (The fact that $C$ is fixed also obviates any analytic issues near $q\approx 1$.) ∎